# EC-Diffuser: Multi-Object Manipulation via Entity-Centric Behavior Generation

### v2

@@ 1 Introduction
Multi-object environments pose significant challenges for autonomous agents due to the combinatorial complexity of both the state and goal spaces as well as of the desired behaviors. Assuming an $n$-object environment with $m$ possible single-object goal configurations, there are $m^{n}$ total goal configurations and $n!$ orderings of the objects to manipulate in sequence. When learning from offline data, one cannot expect an agent to encounter all possible combinations of objects and desired tasks during training due to either time/compute constraints or lack of such data. We therefore require that our agent generalize to novel compositions of objects and/or tasks it has seen during training, i.e. require compositional generalization (Lin et al., 2023).
@@ 2 Related Work
Multi-object Manipulation from Pixels: Previous work using single-vector representations of image observations for control (Levine et al., 2016; Nair et al., 2018; Hafner et al., 2023; Lee et al., 2024) fall short compared to methods that leverage object-centric representations (Zadaianchuk et al., 2021; Yoon et al., 2023; Haramati et al., 2024; Ferraro et al., 2023; Zhu et al., 2022; Shi et al., 2024) in multi-object environments. Several works have studied compositional generalization in this setting. Collaboration et al. (2023); Du et al. (2023; 2024); Zhu et al. (2024) learn directly from pixel observations with large-scale networks, data and compute and rely on scale for possible generalization and transfer. Other works have proposed approaches that account for the underlying combinatorial structure in object manipulation environments to achieve more systematic compositional generalization (Zhao et al., 2022; Chang et al., 2023; Haramati et al., 2024), requiring significantly less data and compute. We continue this line of work, dealing with the setting of learning from demonstrations and the various distinct challenges it introduces.
@@ 4.1 Object-centric Representation with DLP
We first extract a compact, object-centric representation from pixel observations. Given image observations of state $s$ [^2] We note that the true state of the environment $s$ is not observed, and the learned DLP representation is trained purely from pixels. from $V$ different viewpoints, $(I^{0}_{s},...,I^{V-1}_{s})$, we encode each image with a DLPv2 (Daniel & Tamar, 2024) encoder $\phi_{\text{DLP}}$. The resulting representation, as described in Section 3, is a set of $M$ latent particles for each view $v$, denoted by $Z^{v}_{s}=\{z^{v,i}_{s}\}^{M-1}_{i=0}$, where $Z^{v}_{s}=\phi_{\text{DLP}}(I^{v}_{s})$. It is important to note that there is no correspondence between particles from different views (i.e. $z^{v^{\prime},i}_{s}$ and $z^{v^{\prime\prime},i}_{s}$ can represent different objects), nor is there correspondence between particles from different states (i.e. $z^{v,i}_{s^{\prime}}$ and $z^{v,i}_{s^{\prime\prime}}$ can represent different objects). These properties of the DLP representation require a permutation-equivariant policy network architecture, which we describe in the following section.
@@ 4.3 Entity-Centric Diffuser for Goal-Conditioned Behavioral Cloning
We adapt EC-Diffuser to GCBC tasks by incorporating conditioning variables into the diffusion process. Formally, the diffusion process operates over future states and actions: $\mathbf{x}_{0}=\{(Z^{v}_{\tau},a_{\tau})\}_{v=0,\ldots,{V-1},\tau=1,\ldots,H-1}$, with the current state and goal serving as conditional variables: $\mathbf{c}_{g}=\{(Z^{v}_{0},Z^{v}_{g})\}_{v=0,\ldots,{V-1}}$. We define the goal as the last timestep in the demonstration trajectory, i.e., $Z^{v}_{g}=Z^{k}_{T}$, where $T$ is the trajectory length. To train EC-Diffuser, we normalize all input features (DLP’s features and actions) to $[-1,1]$ and employ an $l_{1}$ loss on both states and actions. One might question the effectiveness of using $l_{2}$ or $l_{1}$ losses directly on unordered set inputs. Typically, generating unordered sets like point clouds calls for set-based metrics such as Chamfer distance to compare set similarity. However, in our case, the objective is particle-wise denoising: noise is added independently to each particle, and the denoising process neither imposes nor requires any specific ordering of set elements. Furthermore, we leverage the Transformer’s permutation-equivariant structure by omitting positional embeddings within the set of particles. These factors enable the simple $l_{1}$ loss function to work effectively with diffusion, aligning with previous works that applied diffusion to point clouds (Vahdat et al., 2022; Melas-Kyriazi et al., 2023).
@@ 5.2 Compositional Generalization
Figure 4: PushCube Generalization – Agents trained on $3$ objects and evaluated on $1-6$ objects. See Table 2 for quantitative results. (a) Visualization of a PushCube generalization task. (b), (c) Histograms of per-object goal-reaching success ($n$ out of $6$) in the PushCube generalization task with $6$ objects.
Number of Objects $1$ $2$ $3$ (training) $4$ $5$ $6$
@@
Table 2: PushCube generalization results. The success fraction (number of successful objects / total number of objects) is reported for different numbers of cubes. Agents are trained on $3$ objects with colors randomly selected from $6$ options. Higher values correspond to better performance The best values are in bold.
@@
Figure 5: PushT Generalization – agents trained on $3$ objects with $3$ goals and are tested in different combinations of objects and goals. See Table 3 for quantitative results. (a) Visualization of a generalization task from $3$ to $4$ objects. (b), (c) Histograms of orientation difference (in radians) in the generalization task depicted in (a).
# of Objects / # of Goals $3$/$1$ $3$/$2$ $3$/$3$ (training) $4$/$1$ $4$/$2$ $4$/$3$
@@
Method $1$ Cube $2$ Cubes $3$ Cubes
@@ 5.3 Ablation Studies
Figure 7: Visualization of generated DLP states. The DLP states produced by EC-Diffuser are decoded into images using the pre-trained DLP decoder. (Top) Generated frames for the PushCube generalization task: the model is trained with $3$ cubes and can generate trajectories involving $5$ cubes. (Bottom) Generated frames for the PushT generalization task: the model is trained with $3$ T-blocks and can generate sequences containing $4$. The objects of interest and their respective goals in the generated frames are highlighted with an orange circle.
@@ 5.4 Discussion on State Generation
Additionally, we analyze how well the model maintains the temporal consistency of particles across frames. To do this, we first label the particles according to the object they represent at each timestep in a trajectory. We then perform a T-test comparing the attention values between particles representing the same object over time against the attention values between two random particles. The results show that particles representing the same object over time have significantly higher attention values, with a T-test yielding a p-value of $0.058.367\mathrm{e}{-6}<$ 0.05. This demonstrates that the model implicitly matches objects and enforces object consistency over time, aiding in predicting multi-object dynamics.
@@ Appendix B Environments and Datasets
PushCube: IsaacGym-based (Makoviychuk et al., 2021) tabletop manipulation environment introduced in  Haramati et al. (2024). A Franka Panda robotic arm is required to push cubes in different colors to goal positions specified by images. The agent perceives the environment from two views (front and side) and performs actions in the form of deltas in the end effector coordinates $a=\left(\Delta x_{ee},\Delta y_{ee},\Delta z_{ee}\right)$. Demonstration data for each task (number of objects) was collected by deploying an ECRL (Haramati et al., 2024) state-based agent trained on the corresponding number of objects. We collect $2000$ trajectories per task, each containing $30$, $50$, $100$ transitions for $1$, $2$, $3$ objects respectively.
@@
PushT: IsaacGym-based (Makoviychuk et al., 2021) tabletop manipulation environment introduced in  Haramati et al. (2024). A Franka Panda robotic arm is required to push T-shaped blocks to goal orientations specified by images. The object position is not considered part of the task in this setting. The agent perceives the environment from two views (front and side) and performs actions in the form of deltas in the end effector coordinates $a=\left(\Delta x_{ee},\Delta y_{ee},\Delta z_{ee}\right)$. Demonstration data for each task (number of objects) was collected by deploying an ECRL (Haramati et al., 2024) state-based agent trained on the corresponding number of objects. We collect $2000$ trajectories per task, each containing $50$, $100$, $150$ transitions for $1$, $2$, $3$ objects respectively.
@@ D.1 Additional Metrics
Table 11: Our method’s compositional generalization performance on different numbers of cubes in the PushCube environment, trained on 3 objects in random colors chosen out of $6$ options. The values are computed as the mean over $96$ randomly initialized configurations, and standard deviations are calculated across 5 seeds.
@@ D.2 Additional Object-centric Representations
Input Variation $1$ Cube $2$ Cubes $3$ Cubes
@@ D.3 Generalization to Unseen Objects
Cube Variation $1$ Cube $2$ Cubes $3$ Cubes

### v3

@@ 1 Introduction
Multi-object environments pose significant challenges for autonomous agents due to the combinatorial complexity of both the state and goal spaces as well as of the desired behaviors. Assuming an $n$-object environment with $m$ possible single-object goal configurations, there are $m^{n}$ total goal configurations and $n!$ orderings of the objects to manipulate in sequence. When learning from offline data, one cannot expect an agent to encounter all possible combinations of objects and desired tasks during training due to either time/compute constraints or lack of such data. We therefore require that our agent generalize to novel compositions of objects and/or tasks it has seen during training, i.e. require compositional generalization (Lin et al., 2023a).
@@ 2 Related Work
Multi-object Manipulation from Pixels: Previous work using single-vector representations of image observations for control (Levine et al., 2016; Nair et al., 2018; Hafner et al., 2023; Lee et al., 2024) fall short compared to methods that e object-centric representations (Zadaianchuk et al., 2021; Yoon et al., 2023; Haramati et al., 2024; Ferraro et al., 2023; Zhu et al., 2022; Shi et al., 2024; Lin et al., 2023b) in multi-object environments. Several works have studied compositional generalization in this setting. Collaboration et al. (2023); Du et al. (2023; 2024); Zhu et al. (2024) learn directly from pixel observations with large-scale networks, data and compute and rely on scale for possible generalization and transfer. Other works have proposed approaches that account for the underlying combinatorial structure in object manipulation environments to achieve more systematic compositional generalization (Zhao et al., 2022; Chang et al., 2023; Haramati et al., 2024), requiring significantly less data and compute. We continue this line of work, dealing with the setting of learning from demonstrations and the various distinct challenges it introduces.
@@ 4.1 Object-centric Representation with DLP
We first extract a compact, object-centric representation from pixel observations. Given image observations of state ss [^2] We note that the true state of the environment $s$ is not observed, and the learned DLP representation is trained purely from pixels. from $V$ different viewpoints, $(I^{0}_{s},...,I^{V-1}_{s})$, we encode each image with a DLPv2 (Daniel & Tamar, 2024) encoder $\phi_{\text{DLP}}$. The resulting representation, as described in Section 3, is a set of $M$ latent particles for each view $v$, denoted by $Z^{v}_{s}=\{z^{v,i}_{s}\}^{M-1}_{i=0}$, where $Z^{v}_{s}=\phi_{\text{DLP}}(I^{v}_{s})$. It is important to note that there is no correspondence between particles from different views (i.e. $z^{v^{\prime},i}_{s}$ and $z^{v^{\prime\prime},i}_{s}$ can represent different objects), nor is there correspondence between particles from different states (i.e. $z^{v,i}_{s^{\prime}}$ and $z^{v,i}_{s^{\prime\prime}}$ can represent different objects). These properties of the DLP representation require a permutation-equivariant policy network architecture, which we describe in the following section.
@@ 4.3 Entity-Centric Diffuser for Goal-Conditioned Behavioral Cloning
We adapt EC-Diffuser to GCBC tasks by incorporating conditioning variables into the diffusion process. Formally, the diffusion process operates over future states and actions: $\mathbf{x}_{0}=\{(Z^{v}_{\tau},a_{\tau})\}_{v=0,\ldots,{V-1},\tau=1,\ldots,H-1}$, with the current state and goal serving as conditional variables: $\mathbf{c}_{g}=\{(Z^{v}_{0},Z^{v}_{g})\}_{v=0,\ldots,{V-1}}$. We define the goal as the last timestep in the demonstration trajectory, i.e., $Z^{v}_{g}=Z^{k}_{T}$, where $T$ is the trajectory length. To train EC-Diffuser, we normalize all input features (DLP’s features and actions) to [1,1][-1,1] and employ an $l_{1}$ loss on both states and actions. One might question the effectiveness of using $l_{2}$ or $l_{1}$ losses directly on unordered set inputs. Typically, generating unordered sets like point clouds calls for set-based metrics such as Chamfer distance to compare set similarity. However, in our case, the objective is particle-wise denoising: noise is added independently to each particle, and the denoising process neither imposes nor requires any specific ordering of set elements. Furthermore, we leverage the Transformer’s permutation-equivariant structure by omitting positional embeddings within the set of particles. These factors enable the simple $l_{1}$ loss function to work effectively with diffusion, aligning with previous works that applied diffusion to point clouds (Vahdat et al., 2022; Melas-Kyriazi et al., 2023).
@@ 5.2 Compositional Generalization
Figure 4: PushCube Generalization – Agents trained on 33 objects and evaluated on 161-6 objects. See Table 2 for quantitative results. (a) Visualization of a PushCube generalization task. (b), (c) Histograms of per-object goal-reaching success ($n$ out of 66) in the PushCube generalization task with 66 objects.
Number of Objects 11 22 33 (training) 44 55 66
@@
Table 2: PushCube generalization results. The success fraction (number of successful objects / total number of objects) is reported for different numbers of cubes. Agents are trained on 33 objects with colors randomly selected from $6$ options. Higher values correspond to better performance The best values are in bold.
@@
Figure 5: PushT Generalization – agents trained on 33 objects with 33 goals and are tested in different combinations of objects and goals. See Table 3 for quantitative results. (a) Visualization of a generalization task from 33 to $4$ objects. (b), (c) Histograms of orientation difference (in radians) in the generalization task depicted in (a).
# of Objects / # of Goals 33/11 33/22 33/33 (training) 44/11 44/22 44/33
@@
Method 11 Cube 22 Cubes 33 Cubes
@@ 5.3 Ablation Studies
Figure 7: Visualization of generated DLP states. The DLP states produced by EC-Diffuser are decoded into images using the pre-trained DLP decoder. (Top) Generated frames for the PushCube generalization task: the model is trained with $3$ cubes and can generate trajectories involving 55 cubes. (Bottom) Generated frames for the PushT generalization task: the model is trained with $3$ T-blocks and can generate sequences containing 44. The objects of interest and their respective goals in the generated frames are highlighted with an orange circle.
@@ 5.4 Discussion on State Generation
Additionally, we analyze how well the model maintains the temporal consistency of particles across frames. To do this, we first label the particles according to the object they represent at each timestep in a trajectory. We then perform a T-test comparing the attention values between particles representing the same object over time against the attention values between two random particles. The results show that particles representing the same object over time have significantly higher attention values, with a T-test yielding a p-value of $8.367\mathrm{e}{-6}<0.05$. This demonstrates that the model implicitly matches objects and enforces object consistency over time, aiding in predicting multi-object dynamics.
@@ Appendix B Environments and Datasets
PushCube: IsaacGym-based (Makoviychuk et al., 2021) tabletop manipulation environment introduced in  Haramati et al. (2024). A Franka Panda robotic arm is required to push cubes in different colors to goal positions specified by images. The agent perceives the environment from two views (front and side) and performs actions in the form of deltas in the end effector coordinates $a=\left(\Delta x_{ee},\Delta y_{ee},\Delta z_{ee}\right)$. Demonstration data for each task (number of objects) was collected by deploying an ECRL (Haramati et al., 2024) state-based agent trained on the corresponding number of objects. We collect $2000$ trajectories per task, each containing $30$, $50$, $100$ transitions for $1$, 22, 33 objects respectively.
@@
PushT: IsaacGym-based (Makoviychuk et al., 2021) tabletop manipulation environment introduced in  Haramati et al. (2024). A Franka Panda robotic arm is required to push T-shaped blocks to goal orientations specified by images. The object position is not considered part of the task in this setting. The agent perceives the environment from two views (front and side) and performs actions in the form of deltas in the end effector coordinates $a=\left(\Delta x_{ee},\Delta y_{ee},\Delta z_{ee}\right)$. Demonstration data for each task (number of objects) was collected by deploying an ECRL (Haramati et al., 2024) state-based agent trained on the corresponding number of objects. We collect $2000$ trajectories per task, each containing $50$, $100$, $150$ transitions for $1$, 22, 33 objects respectively.
@@ D.1 Additional Metrics
Table 11: Our method’s compositional generalization performance on different numbers of cubes in the PushCube environment, trained on 3 objects in random colors chosen out of 66 options. The values are computed as the mean over $96$ randomly initialized configurations, and standard deviations are calculated across 5 seeds.
@@ D.2 Additional Object-centric Representations
Input Variation 11 Cube 22 Cubes 33 Cubes
@@ D.3 Generalization to Unseen Objects
Cube Variation 11 Cube 22 Cubes 33 Cubes